OpenStudy (sleepyjess):

Find the sum of the first 12 terms of the sequence. 1, -4, -9, -14, . . .

OpenStudy (sleepyjess):

@hartnn

hartnn (hartnn):

do u see any pattern in that sequence?

OpenStudy (sleepyjess):

subtracting 5

hartnn (hartnn):

correct. thats your "common difference" = d d = 2nd term - 1st term = 3rd term - 2nd term = ... and so on

hartnn (hartnn):

now use the basic sum of arithmetic series formula! \(\Large S_n = (n/2) (2a+(n-1)d)\) a= 1st term and d = -5 and n = 12 = number of terms!

OpenStudy (sleepyjess):

\(\sf S_n = (n/2)(2a+(n-1)d)\) \(\sf S_n = (12/2)(2(1)+(12-1)-5)\) \(\sf S_n = 6(2+11-5)\) \(\sf S_n = 6*8\) \(\sf S_n = 48\)

hartnn (hartnn):

its (n-1)d so \((12-1)(-5) = 11 \times (-5) = -55\)

OpenStudy (sleepyjess):

oh, I didn't realize that was multiplying, not subtracting :/

hartnn (hartnn):

try again? :)

OpenStudy (sleepyjess):

\(\sf S_n = (12/2)(2(1)+(12-1)-5)\) \(\sf S_n = (6)(2+(11)-5)\) \(\sf S_n = (6)(2-55)\) \(\sf S_n = (6)(-53)\) \(\sf S_{12} = -318\)

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (sleepyjess):

Yay!

hartnn (hartnn):

go through the link of the tutorial once, that i pmed you :)

OpenStudy (sleepyjess):

:) Thank you